OpenStudy (anonymous):

solve the following equation showing your work. check your answer and make sure list any restrictions on x. x/(x+1)=1/3+2/(x-2)

OpenStudy (anonymous):

kinda confusing, found the common multiplus, you know the restriction, restriction are simple, denominaor cannot be 0

OpenStudy (anonymous):

is it? x+1? restriction is -1

OpenStudy (anonymous):

ok, u should get the quadratic equation 2x^2 - 7x - 4 = o, solve from there

OpenStudy (anonymous):

if any of the solutions include, x = -1 or x = 2, reject them!

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

x=-1/2 x=4 x cannot = -1 or 2 ??

OpenStudy (anonymous):

those r your answers, yep

OpenStudy (anonymous):

good job!

OpenStudy (anonymous):

yayy thnx!